OpenStudy (anonymous):

the point P(a, klogea) lies on the curve y =klogex where "a" and "k" are constants (c) for what values of "a" does the tangent T cross the x- axis to the LEFT of x=0

OpenStudy (anonymous):

tangent T : kx-ay+ak(ln^2-1)

OpenStudy (anonymous):

go get 'em i am retiring. \[\color{green}{\text{good night!}}\]

myininaya (myininaya):

\[y=\frac{k}{a}*x+k*lna-k\] I believe this was the tangent line to find where the tangent line crosses the x-axis sey y=0 solve for x \[0=\frac{k}{a}*x+k*lna-k\] \[\frac{-k}{a}*x=k*lna-k\] \[x=\frac{-a}{k}*(k*lna-k)=-a(lna-1)\]

myininaya (myininaya):

\[(-a(lna-1),0)\] is the tangent line's x-intercept bye satellite

OpenStudy (anonymous):

sleep tight@ satellite73 and thanks for the help"mate"!

OpenStudy (anonymous):

so when does it cross the x axis to the left of x=0?